A circular current loop of radius 6 angstrom and a current of I pico Ampere is placed in a magnetic field of strength B T. find the potential energy of the magnetic dipole.
I = 69.20 B = 13.40
1
Expert's answer
2020-05-25T10:56:58-0400

Given data

Magnetic field is "B=13.40 T"

Current "I=69.20*10^{-12} A"

Radius "R=6*10^{-10} m"

The potential energy of the magnetic dipole is

"U=-IAB=-I\pi R^2B"

"=-(69.20*10^{-12})\pi (6*10^{-10})^2(13.40)
=-1.05*10^{-27} J"

Here negative sign indicates the dipole moment is parallel to magnetic field.

When the dipole moment is anti parallel to magnetic field, then the potential energy of the magnetic dipole moment is positive value.